OpenStudy (anonymous):

The sum of three different negative integers equals -9. What is the least value that one of these integers could have?

OpenStudy (anonymous):

-8 -6 -5 -4 -1

OpenStudy (zzr0ck3r):

-7 is not on there? -7+(-1)+(-1)

OpenStudy (anonymous):

nope :(

OpenStudy (anonymous):

(-4)+(-2)+(-3)=-9? so -4?

OpenStudy (zzr0ck3r):

-6+(-2)+(-1) = -9 -6<-4

OpenStudy (zzr0ck3r):

but still -7 I think is the answer the smallest negative integer is -1 -1+-1+x=-9 so x=-7

OpenStudy (anonymous):

-7 isn't any of the answer choices, so it's -6 for this question

OpenStudy (zzr0ck3r):

err the closest to 0 is -1 not smallest...

OpenStudy (johnweldon1993):

@zzr0ck3r "three DIFFERENT negative integers" :)

OpenStudy (zzr0ck3r):

there we go:)

OpenStudy (zzr0ck3r):

helps to read, yes -6

OpenStudy (zzr0ck3r):

ty @johnweldon1993

OpenStudy (johnweldon1993):

lol no problem!

OpenStudy (anonymous):

My point that -7 isn't in the answer choices, so the answer for the question is -6, which is B.

OpenStudy (johnweldon1993):

Yes @sakigirl the answer is -6

OpenStudy (zzr0ck3r):

yes, im wrong about -7 because I said -7+-1+-1 =-9, but -1=-1 and the question said "different" so, its not that -6 is the best choice in the list, its that -6 is the best choice period.

OpenStudy (anonymous):

ok cool bro w/e u say

OpenStudy (zzr0ck3r):

lol